OpenStudy (anonymous):

Polynomial problem is: degree 3 polynomial that goes through (5, -12.6) find the polynomial. Any ideas?

OpenStudy (goformit100):



OpenStudy (anonymous):

There are only so many possible answers to this :D

OpenStudy (anonymous):

Did that an the answer I got the math system says is wrong. My answer was: f(x) = x^3(x+2)^2(x-2)

OpenStudy (anonymous):

Thing is, when you let x = 5, f(x) has to be equal to -12.6

OpenStudy (mertsj):

f(x)=ax^3 -12.6=a(5)^3 a=-.1008 f(x=-.1008x^3

OpenStudy (anonymous):

\[\large f(x)=-0.1x^3-0.1=-\frac{x^3+1}{10}\]

OpenStudy (anonymous):

According to WEBwork that is not the correct answer.

OpenStudy (mertsj):

It is probably a simple matter of the form. What do the directions say about entering answers?

OpenStudy (anonymous):

The = sign cannot be used as it is an unacceptable character

OpenStudy (mertsj):

No other directions? How does it say to enter fractions?

OpenStudy (anonymous):

use the division sign for fractions /